# Collision Avoidance and Geofencing for Fixed-wing Aircraft with Control Barrier Functions

### v2

@@ IV-A Position-based CBF Candidates
where $h_{\rm p}$ is a position-based CBF candidate described below. We seek to ensure ${(\mathbf{r}(t),t)\in\mathcal{S}}$, ${\forall t\geq 0}$ for all ${(\mathbf{r}_{0},0)\in\mathcal{S}}$.
@@ IV-A3 Simultaneous collision avoidance and geofencing
Given $N_{\rm c}$ number of safety constraints, these can be combined into a single CBF candidate $h_{\rm p}(\mathbf{r}(t),t)$ based on formulas (27) and (31). For simplicity, we assume that safety constraints are linked with AND logic (i.e., avoiding collision with aircraft 1 AND aircraft 2, etc., AND obeying geofence 1 AND geofence 2, and so on). This leads to the CBF candidate:
@@
with ${\kappa>0}$. Note that ${h_{\rm p}(\mathbf{r},t)=h_{{\rm p},i}(\mathbf{r},t)}$ if ${N_{\rm c}=1}$.
@@
is independent of the controller and may not always hold. As a result, (18) in Definition 2 is not satisfied, $h_{\rm p}$ is not a CBF, and $h_{\rm p}$ cannot be used directly to synthesize safe controllers for (11). We rather call $h_{\rm p}$ as CBF candidate, and below we use it to construct a valid CBF $h$ for controller synthesis.
@@ IV-B Run-time Assurance with Velocity-based Extended CBF
[…]
@@
First, we employ so-called extended CBFs (or higher-order CBFs) proposed in [28, 29]. Extended CBFs, defined as follows, depend on the velocity $\mathbf{v}$:
@@
cf. (36). For simultaneous collision avoidance and geofencing, the individual extended CBFs can be combined as in (27):
@@
Importantly, enforcing ${h(\mathbf{x},t)\geq 0}$ implies that (39) holds with ${\alpha_{\rm p}(r)=\gamma_{\rm p}r}$. Therefore, based on applying Theorem 1 twice, the safety filter (23) guarantees ${h(\mathbf{x}(t),t)\geq 0}$, ${t\geq 0}$, that ultimately results in ${h_{\rm p}(\mathbf{r}(t),t)\geq 0}$, ${\forall t\geq 0}$ (if ${h(\mathbf{x}_{0},0)\geq 0}$ and ${h_{\rm p}(\mathbf{r}_{0},0)\geq 0}$ hold). This means guaranteed safety in collision avoidance and geofencing.
@@ IV-C Run-time Assurance with Backstepping-based CBF
While the extended CBF-based RTA does not leverage all possible behaviors (acceleration, rolling, pitching and yawing), this can be overcome by other CBF choices [30]. In particular, the method of backstepping [31] offers a systematic procedure to synthesize CBFs for cascaded systems like (5). We proceed with backstepping to design a CBF in two steps:
@@
$\begin{split}\mathbf{a}_{\rm s}(\mathbf{x},t)&=\Lambda\big{(}a_{\rm e}(\mathbf{x},t),\|\mathbf{b}_{\rm e}(\mathbf{x},t)\|\big{)}\mathbf{W}_{\rm e}\mathbf{b}_{\rm e}(\mathbf{x},t)^{\top},\\ a_{\rm e}(\mathbf{x},t)&=\dot{h}_{\rm e}\big{(}\mathbf{r},\mathbf{v}(\boldsymbol{\zeta}),t,\mathbf{0}\big{)}+\alpha_{\rm e}\big{(}h_{\rm e}\big{(}\mathbf{r},\mathbf{v}(\boldsymbol{\zeta}),t\big{)}\big{)},\\ \mathbf{b}_{\rm e}(\mathbf{x},t)&=\frac{\partial h_{\rm e}}{\partial\mathbf{v}}\big{(}\mathbf{r},\mathbf{v}(\boldsymbol{\zeta}),t\big{)}\mathbf{W}_{\rm e}.\end{split}$ (46)
@@
The backstepping-based CBF can be established for collision avoidance (with $h_{{\rm e},i}$ from (41)), geofencing (with $h_{{\rm e},i}$ from (42)), and the combination thereof (with $h_{\rm e}$ from (43)). Since the backstepping calculations are nontrivial, it is advantageous to first combine multiple safety constraints into a single one before executing backstepping rather than doing backstepping multiple times and then combining. The safety filter (23) yields safety w.r.t. the backstepping-based CBF, ${h_{\rm b}(\mathbf{x}(t),t)\geq 0}$, ${\forall t\geq 0}$, which implies safety w.r.t. the extended CBF, ${h_{\rm e}\big{(}\mathbf{r}(t),\mathbf{v}(\boldsymbol{\zeta}(t)),t\big{)}\geq 0}$, ${\forall t\geq 0}$, which finally leads to safe behavior considering the position-based CBF candidate, ${h_{\rm p}(\mathbf{r}(t),t)\geq 0}$, ${\forall t\geq 0}$.
The behavior of the aircraft with the proposed safety-critical controller is showcased in Fig. 5. Model (11) is simulated with the safety filter (23)-(24), the backstepping-based CBF (48), and the parameters in Table I. The simulated scenario is simultaneous collision avoidance and geofencing where the geofence consists of two planar boundaries. The aircraft’s RTA system with the proposed backstepping-based CBF guarantees safety with expected behavior. First, the safety filter intervenes and makes the aircraft accelerate, pitch up and turn left to avoid collision with the other aircraft. Then, the aircraft is forced to turn right to avoid crossing the two geofence boundaries. In this case the safety filter never stops intervening, as the aircraft keeps moving parallel to the geofence rather than returning to the original desired trajectory. Throughout the motion, the backstepping-based CBF $h_{\rm b}$ is kept nonnegative, which results in maintaining the three position-based CBFs $h_{{\rm p},i}$ (and their smooth under-approximation $h_{\rm p}$) nonnegative too, as highlighted by Fig. 5(b). These indicate that the underlying maneuvers are executed with guaranteed safety.
@@ IV-D Model-free Run-time Assurance
The main challenge of designing an RTA system is the synthesis of a CBF that respects the cascaded structure of the underlying dynamics. This difficulty would vanish if we could use the position-based CBF candidate $h_{\rm p}$ directly in a model-free fashion. Next, we discuss a simple model-free RTA idea, originally introduced in [32] for robotic systems, and describe its benefits and drawbacks relative to the cascaded model-based RTA presented above.
The proposed model-free approach relies on the ability of the aircraft to track a commanded velocity. Velocity tracking (or trajectory tracking in general) is well-established for many aircraft, for example, the desired controller $\mathbf{k}_{\rm d}$ itself may often be a tracking controller. Such tracking controllers solve a stability problem (associated with Lyapunov theory) that is closely related to ensuring safety (associated with CBF theory) [32]. As a matter of fact, the tracking controller $\mathbf{k}_{\rm d}$ in Appendix B is constructed through a backstepping procedure like the safety filter above. Thus, we intend to leverage the tracking controller (that already accounts for the cascaded dynamics) to address safety in a model-free fashion without going through a complicated control design procedure again.
@@
Equation (50) represents a model-free RTA system which does not involve information about the model (11) but only relies on the expressions (33)-(38) of the position-based CBF candidate $h_{\rm p}$. Tracking the safe velocity $\mathbf{v}_{\rm s}(\mathbf{r},t)$ yields safe behavior under certain assumptions about the tracking controller, as established below. Note that making $\mathbf{v}_{\rm s}$ to be smooth with the smooth safety filter formula facilitates velocity tracking.
Consider a velocity tracking controller, ${\mathbf{u}=\mathbf{k}(\mathbf{x},t)}$, that establishes exponentially stable tracking of the safe velocity $\mathbf{v}_{\rm s}(\mathbf{r},t)$. That is, consider that there exist a Lyapunov function ${V:\mathbb{R}^{n}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}}$ and ${\lambda>0}$ such that:
@@
where the derivative of $V$ is taken along (11). For example, the tracking controller in Appendix B satisfies this condition. Then, it can be established that the system that tracks $\mathbf{v}_{\rm s}$ is safe if $\alpha_{\rm p}$ is chosen such that ${\alpha_{\rm p}(r)=\gamma_{\rm p}r}$ and ${\gamma_{\rm p}<\lambda}$ hold.
@@
The forward invariance of $\mathcal{S}_{V}$ follows from Theorem 1, by showing that the following inequality holds:
@@
where first we used (54); then we substituted (49) and (53); next we applied the Cauchy-Schwartz inequality and (52); and finally we completed the square. Furthermore, safety w.r.t. $\mathcal{S}_{V}$ implies safety w.r.t. $\mathcal{S}$, ${(\mathbf{x},t)\in\mathcal{S}_{V}\implies(\mathbf{r},t)\in\mathcal{S}}$, since ${h_{\rm p}(\mathbf{r},t)\geq h_{V}(\mathbf{x},t)}$ if ${\gamma_{\rm p}<\lambda}$. ∎
@@
The advantage of this model-free approach is that RTA becomes simpler, and it leverages the existing flight controller (i.e., the desired tracking controller). Furthermore, this approach does not modify the low-level flight control system that keeps the aircraft in the air with desired stable behavior, but only uses high-level commands about which direction and how fast the aircraft should fly. The disadvantage of the model-free approach is that it is hard to tune it in a way that the safe velocity is easy to track. This can be observed in Fig. 6(d)-(e) which highlight that tracking the safe velocity results in larger control inputs than the ones with the model-based RTA in Fig. 5(d)-(e). Of course, incorporating model information enhances the performance of the control design, but at the price of complexity—this is a fundamental trade-off observed in many control systems. Nevertheless, the resulting controllers enjoy the formal safety guarantees provided by CBF theory, with both the model-based and model-free RTA.
@@ IV-E Validation with a Full-Envelope System Identified Model
Finally, to validate that the RTA developed for model (11) could provide safety on a higher-fidelity model, simulations were run using an implementation of the linearized, stitched model of the X-62 VISTA (modified two-seat F-16 aircraft) shown in Fig. 7(a). This is a proprietary dynamical model that was developed based on [35], where flight test data were used to identify and combine (i.e., stitch together) linear state-space models with time delays that represent the six-degrees-of-freedom dynamics at various flight conditions. The VISTA simulation was equipped with a neural network controller (NNC) as the nominal flight controller. The NNC was trained via reinforcement learning [9, 36], and it provides the necessary commands in terms of power lever angle (PLA), roll rate, and normal load factor to achieve stable wing-level flight at given altitude, heading, and speed. With the RTA system proposed in this paper, we seek to modulate the NNC whenever necessary to ensure safety w.r.t. collision avoidance.
@@
The high-fidelity simulation results are shown in Fig. 7. A collision avoidance scenario is considered where the VISTA is initially headed towards another aircraft and an evasive maneuver is required from the RTA. To follow the nominal trajectory, the NNC provides PLA, roll rate, and normal load factor commands which can be converted to desired acceleration and angular velocity (i.e., $A_{\rm T}$, $P$ and $Q$). The RTA modifies these desired values to safe acceleration and angular velocity commands. The safe commands are converted back to PLA, roll rate, and normal load factor, and tracked by the VISTA. The RTA is based on the 3D Dubins model (11) and the backstepping-based CBF (48), which was previously demonstrated in Fig. 5. The corresponding simulated trajectories are shown by panels (b), (c), and (d). Remarkably, the RTA constructed from the lower-fidelity 3D Dubins model is capable of providing the required evasive maneuver, and it successfully guarantees safety even in a high-fidelity simulation—thanks to a carefully designed CBF in (48).

### v3

@@ IV-A Position-based CBF Candidates
where $h_{\rm p}$ is a position-based CBF candidate described below. We seek to ensure ${(\mathbf{r}(t),t)\in\mathcal{S}}$, ${\forall t\geq 0}$ for all ${(\mathbf{r}_{0},0)\in\mathcal{S}}$. Note that while we focus on position constraints only, attitude constraints could be addressed in a similar fashion by constructing CBFs that depend on the Euler angles.
@@ IV-A2 Geofencing
Note that zero is a regular value of $h_{{\rm p},i}$ for both (31) and (34), i.e., the gradient of $h_{{\rm p},i}$ is nonzero when ${h_{{\rm p},i}(\mathbf{r},t)=0}$.
@@ IV-A3 Simultaneous collision avoidance and geofencing
Given $N_{\rm c}$ number of safety constraints, these can be combined into a single CBF candidate $h_{\rm p}(\mathbf{r}(t),t)$ based on formulas (25) and (29). For simplicity, we make the following assumption.
Assumption 1.
Safety constraints are linked with AND logic: one must avoid collision with aircraft 1 AND aircraft 2, etc., AND obey geofence 1 AND geofence 2, and so on.
This assumption is made only to keep the exposition simple, since AND logic only requires (25) for CBF composition. Combination of AND and OR logic could also be handled straightforwardly via combining (25) with (29). Considering Assumption 1 and using (25) leads to the CBF candidate:
@@
with ${\kappa>0}$. Note that ${h_{\rm p}(\mathbf{r},t)=h_{{\rm p},i}(\mathbf{r},t)}$ if ${N_{\rm c}=1}$. We remark that (36) approximates a convex polytope for geofencing when $h_{{\rm p},i}$ are given by (34). Non-convex regions could be obtained as the union of convex polytopes, i.e., by using (29).
@@
is independent of the controller and may not always hold. As a result, (14) in Definition 2 is not satisfied, $h_{\rm p}$ is not a CBF, and $h_{\rm p}$ cannot be used directly to synthesize safe controllers for (7). We rather call $h_{\rm p}$ as CBF candidate, and below we use it to construct a CBF $h$ for controller synthesis.
@@ IV-B Run-time Assurance with Velocity-based Extended CBF
safety filter ${\alpha(r)\!=\!\gamma r}$ ${\gamma\!=\!0.1}$ ${\mathbf{W}\!=\!\begin{bmatrix}6\!\!&\!\!0\!\!&\!\!0\\ 0\!\!&\!\!0.6\!\!&\!\!0\\ 0\!\!&\!\!0\!\!&\!\!0.1\end{bmatrix}}$
@@
First, we employ so-called extended CBFs (or high-order CBFs) proposed in [35, 36]. Extended CBFs, defined as follows, depend on the velocity $\mathbf{v}$:
@@
cf. (34). It can be shown that zero is a regular value of $h_{{\rm e},i}$ in (39)-(40), i.e., the gradient of $h_{{\rm e},i}$ is nonzero when ${h_{{\rm e},i}(\mathbf{r},\mathbf{v},t)=0}$. For simultaneous collision avoidance and geofencing, the extended CBFs can be combined as in (25):
@@
Importantly, enforcing ${h(\mathbf{x},t)\geq 0}$ implies that (37) holds with ${\alpha_{\rm p}(r)=\gamma_{\rm p}r}$. Therefore, based on applying Lemma 2, the safety filter (21) guarantees ${h(\mathbf{x}(t),t)\geq 0}$, ${t\geq 0}$ and ${h_{\rm p}(\mathbf{r}(t),t)\geq 0}$, ${\forall t\geq 0}$ (if ${h(\mathbf{x}_{0},0)\geq 0}$ and ${h_{\rm p}(\mathbf{r}_{0},0)\geq 0}$ hold). This means guaranteed safety in collision avoidance and geofencing. Details and proof of the safety guarantees provided by extended CBFs are in [36].
@@ IV-C Run-time Assurance with Backstepping-based CBF
While the extended CBF-based RTA does not leverage all possible behaviors (acceleration, pitching, rolling, and yawing), this can be done by other CBF choices [47]. In particular, the method of backstepping [37] offers a systematic procedure to synthesize valid CBFs for cascaded systems like (5). Here we use backstepping to construct a CBF for use in RTA rather than construct the control law itself. Importantly, backstepping is able to provide a valid CBF whose derivative is affected by all control inputs, hence the underlying RTA can leverage all aforementioned behaviors for safety—even considering 3D motions and nontrivial cascaded dynamics.
We proceed with backstepping to design a CBF in two steps:
@@
$\begin{split}\mathbf{a}_{\rm s}(\mathbf{x},t)&=\Lambda\big{(}a_{\rm e}(\mathbf{x},t),\|\mathbf{b}_{\rm e}(\mathbf{x},t)\|\big{)}\mathbf{W}_{\rm e}\mathbf{b}_{\rm e}(\mathbf{x},t)^{\top},\\ a_{\rm e}(\mathbf{x},t)&=\dot{h}_{\rm e}\big{(}\mathbf{r},\mathbf{v}(\boldsymbol{\zeta}),t,\mathbf{0}\big{)}+\alpha_{\rm e}\big{(}h_{\rm e}\big{(}\mathbf{r},\mathbf{v}(\boldsymbol{\zeta}),t\big{)}\big{)},\\ \mathbf{b}_{\rm e}(\mathbf{x},t)&=\frac{\partial h_{\rm e}}{\partial\mathbf{v}}\big{(}\mathbf{r},\mathbf{v}(\boldsymbol{\zeta}),t\big{)}\mathbf{W}_{\rm e},\end{split}$ (44)
@@
The backstepping-based CBF can be established for collision avoidance (with $h_{{\rm e},i}$ from (39)), geofencing (with $h_{{\rm e},i}$ from (40)), and the combination thereof (with $h_{\rm e}$ from (41)). Since the backstepping calculations are nontrivial, it is advantageous to first combine multiple safety constraints into a single one before executing backstepping rather than doing backstepping multiple times and then combining. The resulting safety filter (21) is in closed form. It yields safety w.r.t. the backstepping-based CBF, ${h_{\rm b}(\mathbf{x}(t),t)\geq 0}$, ${\forall t\geq 0}$, which implies safety w.r.t. the extended CBF, ${h_{\rm e}\big{(}\mathbf{r}(t),\mathbf{v}(\boldsymbol{\zeta}(t)),t\big{)}\geq 0}$, ${\forall t\geq 0}$, which finally leads to safe behavior considering the position-based CBF candidate, ${h_{\rm p}(\mathbf{r}(t),t)\geq 0}$, ${\forall t\geq 0}$. Note that the backstepping-based CBF (46) contains an additional term compared to the extended CBF (38), which is obtained by adding a term to the position-based CBF candidate. Thus, these CBFs have different numbers of parameters. To compare these CBFs and the resulting controllers, we use the same values for the parameters shared by the various CBFs (see Table I), and the comparison reflects the effect of the added terms.
The behavior of the aircraft with the proposed safety-critical controller is showcased in Fig. 5. Model (7) is simulated with the safety filter (21)-(22), the backstepping-based CBF (46), and the parameters in Table I. The simulated scenario is simultaneous collision avoidance and geofencing where the geofence consists of two planar boundaries. The aircraft’s RTA system with the proposed backstepping-based CBF guarantees safety with expected behavior. First, the safety filter intervenes and makes the aircraft accelerate, pitch up and turn left to avoid collision with the other aircraft. Then, the aircraft is forced to turn right to avoid crossing the two geofence boundaries. In this case the safety filter never stops intervening, as the aircraft keeps moving parallel to the geofence rather than returning to the original desired trajectory. Throughout the motion, the backstepping-based CBF $h_{\rm b}$ is kept nonnegative, which results in maintaining the three position-based CBFs $h_{{\rm p},i}$ (and their smooth under-approximation $h_{\rm p}$) nonnegative too, as highlighted by Fig. 5(b). These indicate that the underlying maneuvers are executed with guaranteed safety. While attitude and input constraints are not enforced by the controller, its parameters were tuned so that Euler angles $\theta$, $\phi$ and inputs $P$, $Q$, $A_{\rm T}$ in Fig. 5 evolve within reasonable limits.
@@ IV-D Model-free Run-time Assurance
The main challenge of designing an RTA system is the synthesis of a CBF that respects the cascaded structure of the underlying dynamics. This difficulty would vanish if we could use the position-based CBF candidate $h_{\rm p}$ directly in a model-free fashion. Next, we discuss a model-free RTA idea, originally introduced in [48] for robotic systems, and describe its benefits and drawbacks relative to the cascaded model-based RTA presented above.
The proposed model-free approach relies on the ability of the aircraft to track a commanded velocity. Velocity tracking (or trajectory tracking in general) is well-established for many aircraft, for example, the desired controller $\mathbf{k}_{\rm d}$ itself may often be a tracking controller. Such tracking controllers solve a stabilization problem (associated with Lyapunov theory) that is closely related to ensuring safety (associated with CBF theory) [48]. As a matter of fact, the tracking controller $\mathbf{k}_{\rm d}$ in Appendix B is constructed through a backstepping procedure like the safety filter above. Thus, we intend to leverage the tracking controller (that already accounts for the cascaded dynamics) to address safety in a model-free fashion without going through a complicated control design procedure again.
@@
Equation (48) represents model-free RTA in the sense that it does not use expressions from model (7) but only the position-based CBF candidate $h_{\rm p}$ in (31)-(36). As opposed, controllers with the extended CBF (38)-(42) and the backstepping-based CBF (46) depended on the model. Tracking the safe velocity $\mathbf{v}_{\rm s}(\mathbf{r},t)$ obtained from the model-free RTA yields safe behavior under certain assumptions about the tracking controller, as established below. Note that making $\mathbf{v}_{\rm s}$ to be smooth with the smooth safety filter formula facilitates velocity tracking.
Assumption 2.
Controller ${\mathbf{u}=\mathbf{k}(\mathbf{x},t)}$ yields exponentially stable tracking of the safe velocity $\mathbf{v}_{\rm s}(\mathbf{r},t)$. That is, there exist a Lyapunov function ${V\!:\mathbb{R}^{n}\!\times\!\mathbb{R}_{\geq 0}\!\to\mathbb{R}_{\geq 0}}$ and ${\lambda\!>\!0}$ such that:
@@
where the derivative of $V$ is taken along (7).
For example, the controller in Appendix B satisfies Assumption 2. Then, it can be established that the system that tracks $\mathbf{v}_{\rm s}$ is safe if $\alpha_{\rm p}$ is chosen such that ${\alpha_{\rm p}(r)=\gamma_{\rm p}r}$ and ${\gamma_{\rm p}<\lambda}$.
@@
The forward invariance of $\mathcal{S}_{V}$ follows from Lemma 1, by showing that the following inequality holds:
@@
$\geq\sigma\bigg{\|}\frac{\partial h_{\rm p}}{\partial\mathbf{r}}(\mathbf{r},t)\bigg{\|}^{2}\!\!\!+\!\dot{h}_{\rm p}\big{(}\mathbf{r},t,\mathbf{v}(\boldsymbol{\zeta})\big{)}\!-\!\dot{h}_{\rm p}\big{(}\mathbf{r},t,\mathbf{v}_{\rm s}(\mathbf{r},t)\big{)}\!+\!\frac{V(\mathbf{x},t)}{2\sigma}$
@@
where first we used (52); second we substituted (47) and (51); then we expressed the difference of the $\dot{h}_{\rm p}$ terms; next we applied the Cauchy-Schwartz inequality and (50); and finally we completed the square. Furthermore, safety w.r.t. $\mathcal{S}_{V}$ implies safety w.r.t. $\mathcal{S}$, ${(\mathbf{x},t)\in\mathcal{S}_{V}\implies(\mathbf{r},t)\in\mathcal{S}}$, since ${h_{\rm p}(\mathbf{r},t)\geq h_{V}(\mathbf{x},t)}$ if ${\gamma_{\rm p}<\lambda}$. ∎
@@
The advantage of this model-free approach is that RTA becomes simpler, and it leverages the existing flight controller (i.e., the desired tracking controller). Furthermore, this approach does not modify the low-level flight control system that flies the aircraft with desired stable behavior, but only uses high-level commands about which direction and how fast the aircraft should fly. The disadvantage of the model-free approach is that it is hard to tune it in a way that the safe velocity is easy to track. This can be observed in Fig. 6(d)-(e) which highlight that tracking the safe velocity results in larger control inputs than the ones with the model-based RTA in Fig. 5(d)-(e).
In summary, the velocity-based extended CBF in Section IV-B may succeed in collision avoidance but may fail geofencing tasks, because it is unable to leverage all control inputs (like roll motion) to guarantee safety. The backstepping-based CBF in Section IV-C and the model-free RTA in Section IV-D both leverage all control inputs and can successfully accomplish simultaneous collision avoidance and geofencing. The model-based RTA with backstepping may be preferable over the model-free RTA in terms of the magnitude of control inputs. Of course, incorporating model information enhances the performance of the control design, but at the price of complexity—this is a fundamental trade-off observed in many control systems. Nevertheless, the resulting controllers enjoy the formal safety guarantees provided by CBF theory, with both the model-based and model-free RTA.
@@ IV-E Validation with a Full-Envelope System Identified Model
Finally, to validate that the RTA developed for model (7) could provide safety on a higher-fidelity model, simulations were run using an implementation of the linearized, stitched model of the X-62 VISTA (modified two-seat F-16 aircraft) shown in Fig. 7(a). This is a proprietary dynamical model that was developed based on [51], where flight test data were used to identify and combine (i.e., stitch together) linear state-space models with time delays that represent the six-degrees-of-freedom dynamics at various flight conditions. The VISTA simulation was equipped with a neural network controller (NNC) as the nominal flight controller. The NNC was trained via reinforcement learning [14, 52], and it provides the necessary commands in terms of power lever angle (PLA), roll rate, and normal load factor to achieve stable wing-level flight at given altitude, heading, and speed. With the RTA system proposed in this paper, we seek to modulate the NNC whenever necessary to ensure safety w.r.t. collision avoidance. This serves as the first step towards validating that the Dubins model-based RTA system can be executed on more realistic models, while a comprehensive validation, including experiments on hardware, is left for future work.
The high-fidelity simulation results are shown in Fig. 7. A collision avoidance scenario is considered where the VISTA is initially headed towards another aircraft and an evasive maneuver is required from the RTA. To track the nominal trajectory, the NNC provides PLA, roll rate, and normal load factor commands which can be converted to desired acceleration and angular velocity (i.e., $A_{\rm T}$, $P$, and $Q$). The RTA modifies these desired values to safe acceleration and angular velocity commands. The safe commands are converted back to PLA, roll rate, and normal load factor, and tracked by the VISTA. The RTA is based on the 3D Dubins model (7) and the backstepping-based CBF (46), which was previously demonstrated in Fig. 5. The corresponding simulated trajectories are shown by Fig. 7(b)-(d). Remarkably, the RTA constructed from the lower-fidelity 3D Dubins model is capable of providing the required evasive maneuver, and it successfully guarantees safety even in a high-fidelity simulation—thanks to a carefully designed CBF in (46).
We remark that the high-fidelity simulation environment was also monitoring additional constraints on the aircraft, such as limits for angle of attack, angle of side slip, pilot $g$ command, roll rate, yaw rate, normal load factor, side load factor, speed, control surface deflections, and their rates. These constraints include some of the states and inputs of the Dubins model, as well as additional quantities that are not present in the model. Upon violation of these constraints, the simulation terminated. By tuning the parameters of the proposed controller, we prevented the violation of these constraints while maintaining safety w.r.t. collision avoidance.
Our future plan is to formally address these state and input constraints by more complex CBF formulations such as the backup set method [15, 28]. This method uses a backup controller that could maintain safe behavior with input-constrained conservative maneuvers like turning the aircraft away from unsafe regions. By forward integrating the closed-loop dynamics and computing the corresponding backup trajectory, safety filters can enforce that this trajectory leads to a backup set without violating safety. This strategy ultimately provides formal guarantees of safety while satisfying input constraints, and its safety filter may integrate a desired flight controller to achieve a higher performance than using a conservative backup controller directly. Developing the required backup controllers and sets, however, needs significant future research.